The following table summarizes a portion of the results for a two-way analysis of variance experiment with no interaction.

Find the missing values in the ANOVA table. (Round intermediate calculations to at least 4 decimal places. Round "MS" to 4 decimal places and "F" to 3 decimal places.)

source of Variation SS df MS F p-value

rows 34.63 2 MSB= factor B= 0.042

columns 146.60 3 MSA= Factor A= 0.003

error 18.56 6 MSE=

total 199.79 11
Answer #1

We are given

SSB = 34.63 , SSA = 146.60 , SSE = 18.56 , SST = 199.79

d.f B = 2 , d.f A = 3 , d.f Error = 6 , d.f total = 11

MSB = SSB /d.f B = 34.63/2 = 17.315

MSA = SSA/ df A = 146.60/3 = 48.8667

MS error = SS error/ df error = 18.56/6 = 3.0933

F​​​​​B = MSB/MS error = 17.315/3.0933 = 5.598

F​​​​​​B = 5.598

F​​​​​A = MSA/ MS error = 48.8667/3.0933 = 15.798

F​​​​​​A = 15.798